A Partial Result on a Problem of Jech

Natasha Dobrinen

In connection with a problem of von Neumann, whether the countable chain condition and the weak (ω, ω)-distributive law characterize measurable algebras among Boolean σ-algebras, I am investigating relationships between distributive laws and Boolean algebraic games. I will present a game-theoretic characterization of the weak (ω, ω)-distributive law in Boolean σ-algebras which partially answers an open question posed by Jech [1].

T. Jech: More game-theoretic properties of Boolean algebras, Ann. Pure and Appl. Logic , 1984, vol 26 pp. 11-29

Address: University of Minnesota - Twin Cities, 206 Church St. SE, 127 Vincent Hall, University of Minnesota, Minneapolis, MN 55414, United States

E-mail: dobrinen@math.umn.edu